A BASELINE SUBTRACTION TOOL

\underline{Z. DUAN} and R. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139; A. F. RUCKSTUHL, Departement P, Zurcher Hochschule Winterthur (Zurich University of Applied Sciences), Postfach 805, 8401 Winterthur, Switzerland; M. JACOBSON, Department of Chemistry, Columbia University, New York, NY 10027; J. DODD, Space Vehicles Directorate, Air Force Research Laboratory, Hanscom Air Force Base, MA 01731.

An Igor macro based on robust baseline estimation is introduced, which uses techniques of robust local regression to estimate baselines in spectra that consist of sharp features superimposed upon a continuous, slowly varying baseline ^{a}.

aA. F. Ruckstuhl, M. P. Jacobson, R. W. Field, and J. A. Dodd, Journal of Quantitative Spectroscopy and Radiative Transfer, 68, 179-193 (2001).